Grade 12Physical Chemistry

Hydrogenation of vegetable ghee at 25 degree celsius reduced pressure of H2 from 2 atm to 1.2 atm in 50 min . The rate of reaction in terms of molarity per second is?

Sai Soumya
8 Years agoGrade 12

1 Answer

Arun
8 Years ago
hydrogenation is a zero order reaction !
the rate of decrease of pressure is given by,
-dp /dt = k(pAa) (pBb )..........assuming that reaction occurs between A and B and (a) and (b) are the orders of the reaction wrt A and B respectively
,,,since this is a zero order reaction, a=b=0
=> k= rate constant =-dp/dt = 0.8 /50 = 0.016
.....hence rate of reaction (in this zero order reaction) =0.016